# Similarity Field Theory: A Mathematical Framework for Intelligence

### v1

@@ Abstract
We posit that persisting and transforming similarity relations form the structural basis of any comprehensible dynamic system. This paper introduces Similarity Field Theory, a mathematical framework that formalizes the principles governing similarity values among entities and their evolution. We define: (1) a similarity field $S:U\times U\to[0,1]$ over a universe of entities $U$, satisfying reflexivity $S(E,E)=1$ and treated as a directed relational field (asymmetry and non-transitivity are allowed); (2) the evolution of a system through a sequence $Z_{p}=(X_{p},\,S^{(p)})$ indexed by $p=0,1,2,\dots$; (3) concepts $K$ as entities that induce fibres $F_{\alpha}(K)=\{E\in U\mid S(E,K)\geq\alpha\}$, i.e., superlevel sets of the unary map $S_{K}(E):=S(E,K)$; and (4) a generative operator $G$ that produces new entities. Within this framework, we formalize a generative definition of intelligence: an operator $G$ is intelligent with respect to a concept $K$ if, given a system containing entities belonging to the fibre of $K$, it generates new entities that also belong to that fibre. Similarity Field Theory thus offers a foundational language for characterizing, comparing, and constructing intelligent systems. We prove two theorems: (i) asymmetry blocks mutual inclusion; and (ii) stability requires either an anchor coordinate or eventual confinement within a level set of $f$. These results ensure that the evolution of similarity fields is both constrained and interpretable, culminating in an exploration of how the framework allows us to interpret large language models and use them as experimental probes into societal cognition.
@@ 1 Introduction
  • A first-principles framework that treats graded, possibly asymmetric similarity as the primitive substrate for cognition and intelligence, with fibres as superlevel sets that define conceptual membership.
  • A sequence-based evolution formalism, including a strengthened stability theorem that characterizes anchors versus eventual confinement and an incompatibility theorem showing how asymmetry forbids reciprocal inclusion under self-set thresholds.
  • An operational, generative definition of intelligence in terms of fibre preservation and coverage/fidelity measures, applicable across models and modalities.
  • A program for interpretability that views networks as compositions of calibrated similarity fields and demonstrates how pairwise typicality probes recover aggregate cognitive structure.
@@ 2.3 Category Theory: A Language of Structure
Category theory provides a language for abstract structures, describing objects and their connections (morphisms). In the standard view, the link between two objects is a discrete, binary matter: a morphism either exists or does not. While powerful, this is insufficient to capture the graded nature of conceptual relations. We use a similarity field $S$ taking values in $[0,1]$ to quantify the strength of belonging/connection; given a concept $K$, we express structure and invariants via threshold sets of $S(\cdot,K)$; generation is “intelligent” iff it preserves and expands these threshold sets along the evolution. We do not read $S$ as a discrete morphism between arbitrary objects; instead, we center on the concept $K$ and study $S(\cdot,K)$ and its fibres $F_{\alpha}(K)$, using strength rather than mere existence to characterize belonging and connection. More importantly, this framework targets any system organized by similarity—from perception and learning to machine models, social bargaining, and institutional design—rather than being confined to categorical settings or specialist uses of “fibre”. Enriched categorical phrasing is only an explanatory lens; our theory and results hold even without adopting it [13, 14].
@@ 2.4.1 Current Explorations and the Need for a Theory
The academic community has made explorations into model interpretability, developing techniques to trace information “circuits” [21], identify “persona vectors” in activation space [25, 26], and perform “representation engineering” [24, 22, 23]. These empirical works enhance our ability to inspect AI systems, but they often lack a unifying mathematical theory that explains how these observed structures emerge from first principles. The Similarity Field Theory framework is proposed not to replace these techniques, but to offer a foundational mathematical perspective, based on how similarity relations constitute concepts, to help unify and guide these interpretability studies at a deeper, structural level.
@@ 2.4.2 Beyond a Purely Statistical View
Viewing a neural network as a statistical learning machine is a dominant and effective practice; it excels at characterizing in-distribution input–output behavior but offers limited insight into the formation of internal conceptual structure and into out-of-distribution (OOD) phenomena. Similarity Field Theory complements this view from the angle of conceptual similarity: we regard a trained model as a complex system composed of nested similarity fields and restate interpretability as a task of conceptual deconstruction—identifying latent concepts formed inside the model and defining them via their fibres (the sets of inputs that elicit high similarity responses). Consequently, the research goal shifts from merely fitting external statistics to understanding and organizing these concept fibres, enabling structured judgments beyond the training distribution in accordance with conceptual similarity rather than being constrained by frequency statistics of the training distribution. For example, if the training data include only “red circles” and “green triangles” and never “green circles,” the model can still classify a green circle as a circle because it matches along the fibre of the concept “circle,” rather than relying on frequency co-occurrence.
@@ 3 Foundational Constructs
A concept, denoted $K$, is an entity that serves as a representation for a class of other entities. It induces a unary map $S_{K}(E):=S(E,K)$ and thereby defines fibres as its superlevel sets.
@@
Definition 5 (Fibre).
Given a concept $K\in U$ and a threshold $\alpha\in[0,1]$, the fibre induced by $K$ is the superlevel set:
@@
Proposition 1 (Fibre monotonicity).
@@
3.1 On the Nature of the Similarity Field
  • Symmetry: We do not assume that $S(E_{1},E_{2})=S(E_{2},E_{1})$. This allows the framework to model asymmetric cognitive phenomena in reality, such as the perceived similarity of a novice to an expert, which is typically higher than the reverse.
  • Transitivity: High similarity need not be transitive, aligning with many cognitive examples.
This deliberate relaxation is motivated by the arguments presented in Section 2, distinguishing Similarity Field Theory from frameworks based on Metric Spaces or strict Equivalence Relations. While a field satisfying all three properties is a valid special case, our broader definition is essential for modeling the continuous and overlapping nature of conceptual structures in the real world.
3.2 Properties of Similarity Fields
@@
    $E_{1}$ belongs to the fibre induced by $E_{2}$ as a concept, with threshold $y$ (i.e., $E_{1}\in F_{y}(E_{2})$).
@@
    $E_{2}$ belongs to the fibre induced by $E_{1}$ as a concept, with threshold $x$ (i.e., $E_{2}\in F_{x}(E_{1})$).
@@
  • From condition 1 being true, by the definition of a fibre $F_{y}(E_{2})=\{E\in U\mid S(E,E_{2})\geq y\}$, we must have $S(E_{1},E_{2})\geq y$. Substituting $x=S(E_{1},E_{2})$, we get $x\geq y$.
  • From condition 2 being true, by the definition of a fibre $F_{x}(E_{1})=\{E\in U\mid S(E,E_{1})\geq x\}$, we must have $S(E_{2},E_{1})\geq x$. Substituting $y=S(E_{2},E_{1})$, we get $y\geq x$.
@@
Significance of Theorem 1: It is precisely the introduction of asymmetry that makes this theorem a fundamental and non-trivial principle. Far from being a mere restatement of definitions, it reveals the inherent limitations and forced imbalance that conceptual membership inevitably faces in an asymmetric world. It establishes the impossibility of reciprocal inclusion: if $E_{1}$’s perspective on $E_{2}$ differs from $E_{2}$’s perspective on $E_{1}$ (i.e., $x\neq y$), then they cannot simultaneously be members of each other’s conceptual fibres when the threshold for inclusion is set by the other’s standard. Thus, this theorem provides a formal explanation for negotiation deadlocks in social interaction: when two parties’ assessments of each other’s positions are asymmetric, and both adopt the other’s standard as the threshold for acceptance, a mutual agreement becomes mathematically impossible unless one or both parties adjust their acceptance criteria (i.e., the fibre threshold).
This principle is illustrated by the Novice ($N$) and Expert ($E$) scenario. Suppose $S(N,E)=0.9$ and $S(E,N)=0.2$. The Novice may belong to the Expert’s fibre ($N\in F_{0.2}(E)$ since $0.9\geq 0.2$), but the theorem guarantees the impossibility of the reverse: the Expert cannot belong to the Novice’s fibre at the Novice’s standard ($E\notin F_{0.9}(N)$ since $0.2<0.9$). The asymmetry in similarity mathematically forbids a symmetric sense of belonging.
@@
Figure 1: The Incompatibility Theorem visualized through a negotiation. Each party’s "acceptance fibre" is defined by the other’s perceived value. Party B’s offer ($S(B,A)=0.3$) is too low to enter Party A’s acceptance fibre (threshold 0.8). Conversely, Party A’s offer ($S(A,B)=0.8$) easily clears Party B’s threshold (0.3). Because the condition is not met for both, mutual agreement is impossible.
@@ 4 System Evolution via Sequences
Eventual confinement and absorbing sets. A subset $\mathcal{A}\subseteq U$ is called an absorbing set for a given evolution sequence if there exists $p^{\star}$ such that for all $p\geq p^{\star}$, the entities under consideration lie in $\mathcal{A}$ and never leave it along the subsequent sequence. When $\mathcal{A}$ is a fibre $F_{\alpha}(K)$, we may say the sequence exhibits eventual confinement inside the fibre of concept $K$.
@@ 5 Generative Dynamics and Intelligence
Within the Similarity Field Theory framework, this principle is expressed formally. The premise, “given entities exemplifying a concept,” corresponds to a system-state $Z_{p}$ that contains a subset of entities $X_{K}\subseteq X_{p}$, where every entity in $X_{K}$ belongs to a fibre $F_{\alpha}(K)$.
Intelligence with respect to a concept $K$ is the capacity of a generative operator $G$ to produce new entities $E^{\prime}\in G(X_{p})$ that also belong to the fibre $F_{\alpha}(K)$ by satisfying the condition:
@@ 5.2 Creativity as Re-Contextualization
Furthermore, our framework provides a formal perspective on creativity. This is not treated as creation ex nihilo, but as the formation of a new concept initiated by the re-contextualization of an existing entity. This occurs when an entity $E$, already a member of a source concept $K_{\text{source}}$, is discovered to also possess high similarity to a different, nascent target concept $K_{\text{target}}$. Prior to this discovery, the fibre of the target concept, $F_{\beta}(K_{\text{target}})$, might have been empty within the system. The act of recognizing that $S(E,K_{\text{target}})\geq\beta$ populates this fibre for the first time, establishing $K_{\text{target}}$ as a valid concept and positioning $E$ as its foundational instance. The history of WD-40 illustrates this. Its formula ($E$) was an entity in the fibre of “aerospace anti-corrosion agent” ($K_{\text{source}}$). The discovery of its utility in household applications was a recognition of its high similarity to the nascent concept of a “general-purpose lubricant” ($K_{\text{target}}$), making it the first entity to populate that conceptual fibre and create a new market. At an abstract level, creativity can also be characterized as re-contextualization: an entity in the fibre of a source concept is recognized as crossing the threshold for a target concept, thereby making the new fibre non-empty for the first time.
@@ 6 Interpretation in Machine Learning
Similarity Field Theory provides a new analytical lens for understanding typical machine learning models. As discussed earlier in our “Beyond a purely statistical view” subsection, our aim here is not to rehash frequency co-occurrence arguments but to treat networks as knowledge structures of concept fibres; we instantiate this perspective with two canonical cases: neural networks and LLMs [6].
6.1 Neural Networks as Composed Similarity Fields
Consider a neural network trained to classify images of a cat. Let $K_{\text{Cat}}$ be the concept “Cat”. The network is trained to approximate a similarity field, $S_{\text{NN}}(E_{\text{input}},K_{\text{Cat}})\to[0,1]$, where the input entity $E_{\text{input}}$ is a vector of pixel values.
Measurement readout.
Internal activations may be unbounded (finite). For interpretation we can fix a strictly increasing bijection $\phi:\mathbb{R}\to(0,1)$ and read similarity as $s=\phi(r)$; when needed we recover the raw scale via $r=\phi^{-1}(s)$. This defines a calibrated readout space used for algebraic/closure statements and does not modify the network’s forward pass or internal value distribution [18].
  1. 1.
    A Single Neuron as a Feature-Similarity Field: A neuron in the first hidden layer receives $E_{\text{input}}$. Via a monotone calibration $\tilde{a}=\phi(a)\in(0,1)$, its activation induces a primitive field $S(E_{\text{input}},K_{\text{feature},\mathbf{w}})$. We set concept self-similarity $S(K_{\text{feature},\mathbf{w}},K_{\text{feature},\mathbf{w}})=1$ and choose a prototype input $E^{\star}$ with $S(E^{\star},K_{\text{feature},\mathbf{w}})=1$ (Fig. 2).
  2. 2.
    The Network as a Learned Composition: Each layer either takes a nonnegative weighted average of calibrated field values, or applies an anchor-preserving aggregator $g_{\ell}:[0,1]^{m}\to[0,1]$ with $g_{\ell}(\mathbf{1})=1$ and coordinatewise monotonicity; hence the layer’s prototype retains value $1$. On calibrated sub-fields, convex-combination and pointwise-product closure keep the range in $[0,1]$ and maintain prototypes at $1$, so the final output remains a valid, high-level similarity field approximating the target $S(E_{\text{input}},K_{\text{Cat}})$ (Fig. 3).
  3. 3.
    Training as Similarity Maximization: The training process optimizes this composed field, seeking to increase the expected similarity for entities that are members of the fibre $F_{\alpha}(K_{\text{Cat}})$.
@@ 6.1 Neural Networks as Composed Similarity Fields
6.2 Large Language Models as Concept Synthesizers
A token is a concept, $K_{\text{token}}$, and an input prompt is a sequence of concepts that are composed to form a single, higher-order concept, $K_{\text{prompt}}$.
The entire trained LLM can be viewed as an entity that represents a vast, complex meta-concept of language, $K_{\text{Language}}$. Simultaneously, it functions as a generative operator, $G$. For a token concept $K_{\text{token}}$, its fibre collects entities in $U$ that align with the token’s usage in human language and culture. When this operator acts upon a prompt, its prediction head compares the synthesized concept $K_{\text{prompt}}$ with every concept in its vocabulary. The resulting probability distribution, after calibration, can be treated as a monotone surrogate of the similarity field values; selecting a high-probability next token $E^{\prime}_{\text{next}}$ thus instantiates $S(E^{\prime}_{\text{next}},K_{\text{prompt}})\geq\alpha$, where $\alpha$ can be a calibrated probability threshold or a rank-based quantile (e.g., top-$k$, top-$p$) [19].
6.3 A Path towards Interpretability
This framework suggests a formal path to achieving model interpretability. Since the entire network is a composition of simpler similarity fields, it can be conceptually deconstructed into its fundamental components. Each neuron computes a field value with respect to a latent concept, $S(E_{\text{input}},K_{\text{latent}})$, so we can understand the model by reverse-engineering these atomic concepts.
The concept represented by a neuron, $K_{\text{latent}}$, can be fully characterized by its fibre: the set of all inputs that cause its calibrated activation to be 1.
@@ 6.3 A Path towards Interpretability
Interpretability then becomes the problem of deconstructing the model into its constituent conceptual fibres and understanding how they are composed.
6.4 As Similarity Inference

### v4

@@ Abstract
We posit that persisting and transforming similarity relations form the structural basis of any comprehensible dynamic system. This paper introduces Similarity Field Theory, a mathematical framework that formalizes the principles governing similarity values among entities and their evolution. We define: (1) a similarity field $S:U\times U\to[0,1]$ over a universe of entities $U$, satisfying reflexivity $S(E,E)=1$ and treated as a directed relational field (asymmetry and non-transitivity are allowed); (2) the evolution of a system through a sequence $Z_{p}=(X_{p},\,S^{(p)})$ indexed by $p=0,1,2,\dots$; (3) concepts $K$ as entities that induce fibers $F_{\alpha}(K)=\{E\in U\mid S(E,K)\geq\alpha\}$, i.e., superlevel sets of the unary map $S_{K}(E):=S(E,K)$; and (4) a generative operator $G$ that produces new entities. Within this framework, we formalize a generative definition of intelligence: an operator $G$ is intelligent with respect to a concept $K$ if, given a system containing entities belonging to the fiber of $K$, it generates new entities that also belong to that fiber. Similarity Field Theory thus offers a foundational language for characterizing, comparing, and constructing intelligent systems. At a high level, this framework reframes intelligence and interpretability as geometric problems on similarity fields—preserving and composing level-set fibers—rather than purely statistical ones. We prove two theorems: (i) asymmetry blocks mutual inclusion; and (ii) stability implies either an anchor coordinate or asymptotic confinement to the target level (up to arbitrarily small tolerance). Together, these results constrain similarity-field evolution and motivate an interpretive lens that can be applied to large language models.
@@ 1 Introduction
  • A first-principles framework that treats graded, possibly asymmetric similarity as the primitive substrate for cognition and intelligence, with fibers as superlevel sets that define conceptual membership.
  • A sequence-based evolution formalism, including a strengthened stability theorem that characterizes anchors versus eventual confinement and an incompatibility theorem showing how asymmetry forbids reciprocal inclusion under cross-referenced thresholds.
  • An operational, generative definition of intelligence in terms of fiber preservation and coverage/fidelity measures.
  • A framework for interpretability that views networks as compositions of calibrated similarity fields.
@@ 2.3 Category Theory: A Language of Structure
Category theory provides a language for abstract structures, describing objects and their connections (morphisms). In the standard view, the link between two objects is a discrete, binary matter: a morphism either exists or does not. While powerful, this is insufficient to capture the graded nature of conceptual relations. We use a similarity field $S$ taking values in $[0,1]$ to quantify the strength of belonging/connection; given a concept $K$, we express structure and invariants via threshold sets of $S(\cdot,K)$; generation is “intelligent” iff it preserves and expands these threshold sets along the evolution. We do not read $S$ as a discrete morphism between arbitrary objects; instead, we center on the concept $K$ and study $S(\cdot,K)$ and its fibers $F_{\alpha}(K)$, using strength rather than mere existence to characterize belonging and connection. More importantly, this framework targets any system organized by similarity—from perception and learning to machine models, social bargaining, and institutional design—rather than being confined to categorical settings or specialist uses of “fiber”. Enriched categorical phrasing is only an explanatory lens; our theory and results hold even without adopting it [10, 8].
@@ 2.4.1 Current Explorations and the Need for a Theory
The academic community has made explorations into model interpretability, developing techniques to trace information “circuits” [12], identify “persona vectors” in activation space [3, 6], and perform “representation engineering” [19, 16, 14]. These empirical works enhance our ability to inspect AI systems, but they often lack a unifying mathematical theory that explains how these observed structures emerge from first principles. The Similarity Field Theory framework is proposed not to replace these techniques, but to offer a foundational, geometric mathematical perspective: it casts models as calibrated similarity fields whose level-set fibers constitute concepts, thereby unifying and guiding interpretability studies at a deeper structural level beyond purely statistical views.
@@ 2.4.2 Beyond a Purely Statistical View
Viewing a neural network as a statistical learning machine is a dominant and effective practice; it excels at characterizing in-distribution input–output behavior but offers limited insight into the formation of internal conceptual structure and into out-of-distribution (OOD) phenomena. Similarity Field Theory complements this view from the angle of conceptual similarity: we regard a trained model as a complex system composed of nested similarity fields and restate interpretability as a task of conceptual deconstruction—identifying latent concepts formed inside the model and defining them via their fibers (the sets of inputs that elicit high similarity responses). Consequently, the research goal shifts from merely fitting external statistics to understanding and organizing these concept fibers, enabling structured judgments beyond the training distribution in accordance with conceptual similarity rather than being constrained by frequency statistics of the training distribution. For example, if the training data include only “red circles” and “green triangles” and never “green circles,” the model can still classify a green circle as a circle because it matches along the fiber of the concept “circle,” rather than relying on frequency co-occurrence.
Box 1  |  Why fiber-based structure can generalize when joint-frequency MLE collapses Consider two attributes—color $C\in\{\mathrm{red},\mathrm{green}\}$ and shape $H\in\{\mathrm{circle},\mathrm{triangle}\}$—and entities of the form $(C,H)$. Suppose the training set contains only $(\mathrm{red},\mathrm{circle})$ and $(\mathrm{green},\mathrm{triangle})$. A joint-frequency maximum-likelihood estimate assigns $\widehat{P}_{\mathrm{MLE}}(C=\mathrm{green},H=\mathrm{circle})=\frac{n(\mathrm{green},\mathrm{circle})}{N}=0,$ so any decision rule that relies on the joint MLE as “support” cannot endorse the unseen combination $(\mathrm{green},\mathrm{circle})$ without introducing extra structure. In Similarity Field Theory (SFT), concepts are modeled as superlevel sets of a similarity field. Let $K$ denote the concept “circle”, and define the fiber $F_{\alpha}(K)=\{E:\ S(E,K)\geq\alpha\}$. Let $X_{K}$ be a small exemplar set intended to lie in the fiber; in the simplest case, $X_{K}=\{x\}$ with $x=(\mathrm{red},\mathrm{circle})$. Define an exemplar-to-concept readout $\hat{S}(E,K)=\max_{x\in X_{K}}\min\big\{S_{K}(E,x),\,S(x,K)\big\},$ where $S_{K}(\cdot,\cdot)$ denotes a $K$-aligned similarity probe (e.g., a shape-focused similarity). Lemma (one-step fiber extension). If there exists $x\in X_{K}$ such that $S(x,K)\geq\alpha$ and $S_{K}(E,x)\geq\alpha$, then $\hat{S}(E,K)\geq\alpha$ and hence $E\in\hat{F}_{\alpha}(K)$. Thus, even when $n(\mathrm{green},\mathrm{circle})=0$, the unseen entity $E=(\mathrm{green},\mathrm{circle})$ can be admitted into the “circle” fiber whenever it is sufficiently similar (in the $K$-aligned sense) to an exemplar already inside the fiber. By contrast, additive smoothing (e.g., Laplace/Dirichlet) redistributes mass to unseen cells without preferentially selecting $(\mathrm{green},\mathrm{circle})$ over other unobserved combinations unless additional factorization or concept-specific structure is imposed.
@@ 3 Foundational Constructs
A concept, denoted $K$, is an entity that serves as a representation for a class of other entities. It induces a unary map $S_{K}(E):=S(E,K)$ and thereby defines fibers as its superlevel sets.
@@
Definition 5 (Fiber).
Given a concept $K\in U$ and a threshold $\alpha\in[0,1]$, the fiber induced by $K$ is the superlevel set:
@@
3.1 On the Nature of the Similarity Field
  • Symmetry: We do not assume that $S(E_{1},E_{2})=S(E_{2},E_{1})$. This allows the framework to model asymmetric cognitive phenomena in reality. Classic work in cognitive psychology has shown that people often judge, for example, “North Korea is similar to China” as more natural, and endorse it more frequently, than the reversed statement “China is similar to North Korea”, even though the underlying pair of countries is the same; such reference-point and salience effects systematically break symmetry in human similarity judgments [17].
  • Transitivity: High similarity need not be transitive. In conceptual space, consider lion (A), domestic cat (B), and household pet (C): $S(A,B)$ is high (both are felids), $S(B,C)$ is high (domestic cats are prototypical pets), yet $S(A,C)$ is much lower (lions are not household pets), yielding high $S(A,B)$ and $S(B,C)$ but comparatively low $S(A,C)$. This pattern departs from metric triangle-inequality assumptions and matches empirical observations about graded category structure and conceptual similarity [17, 5].
This deliberate relaxation is motivated by the arguments presented in the main text, distinguishing Similarity Field Theory from frameworks based on Metric Spaces or strict Equivalence Relations. While a field satisfying all three properties is a valid special case, our broader definition is essential for modeling the continuous and overlapping nature of conceptual structures in the real world.
3.2 Properties of Similarity Fields
Proposition 1 (Fiber monotonicity).
@@ 3.2 Properties of Similarity Fields
    $E_{1}$ belongs to the fiber induced by $E_{2}$ as a concept, with threshold $y$ (i.e., $E_{1}\in F_{y}(E_{2})$).
@@
    $E_{2}$ belongs to the fiber induced by $E_{1}$ as a concept, with threshold $x$ (i.e., $E_{2}\in F_{x}(E_{1})$).
@@
  • From condition 1 being true, by the definition of a fiber $F_{y}(E_{2})=\{E\in U\mid S(E,E_{2})\geq y\}$, we must have $S(E_{1},E_{2})\geq y$. Substituting $x=S(E_{1},E_{2})$, we get $x\geq y$.
  • From condition 2 being true, by the definition of a fiber $F_{x}(E_{1})=\{E\in U\mid S(E,E_{1})\geq x\}$, we must have $S(E_{2},E_{1})\geq x$. Substituting $y=S(E_{2},E_{1})$, we get $y\geq x$.
@@
Significance of Theorem 1: The key point is that asymmetry induces a one-sided constraint: under cross-referenced thresholds, reciprocal inclusion becomes impossible unless the similarity relation is symmetric. This highlights a structural constraint: asymmetry can impose inherent limitations and a forced imbalance on conceptual membership. It establishes the impossibility of reciprocal inclusion: if $E_{1}$’s perspective on $E_{2}$ differs from $E_{2}$’s perspective on $E_{1}$ (i.e., $x\neq y$), then they cannot simultaneously be members of each other’s conceptual fibers when the threshold for inclusion is set by the other’s standard. In a stylized negotiation setting where each party sets its acceptance threshold using the other party’s valuation, the theorem yields a simple infeasibility condition: if the assessments are asymmetric, mutual agreement cannot be achieved under these cross-referenced thresholds unless at least one party adjusts its acceptance criterion, for example by slightly lowering a price (or raising an offer), which effectively shifts the fiber threshold and may, when the parties are near the boundary, turn non-membership into membership.
This principle is illustrated by the Novice ($N$) and Expert ($E$) scenario. Suppose $S(N,E)=0.9$ and $S(E,N)=0.2$. The Novice may belong to the Expert’s fiber ($N\in F_{0.2}(E)$ since $0.9\geq 0.2$), but the theorem guarantees the impossibility of the reverse: the Expert cannot belong to the Novice’s fiber at the Novice’s standard ($E\notin F_{0.9}(N)$ since $0.2<0.9$). The asymmetry in similarity mathematically forbids a symmetric sense of belonging.
@@
Figure 1: The Incompatibility Theorem visualized through a negotiation. Each party’s "acceptance fiber" is defined by the other’s perceived value. Party B’s offer ($S(B,A)=0.3$) is too low to enter Party A’s acceptance fiber (threshold 0.8). Conversely, Party A’s offer ($S(A,B)=0.8$) easily clears Party B’s threshold (0.3). Because the condition is not met for both, mutual agreement is impossible.
@@ 4 System Evolution via Sequences
Eventual confinement and absorbing sets. A subset $\mathcal{A}\subseteq U$ is called an absorbing set for a given evolution sequence if there exists $p^{\star}$ such that for all $p\geq p^{\star}$, the entities under consideration lie in $\mathcal{A}$ and never leave it along the subsequent sequence. When $\mathcal{A}$ is a fiber $F_{\alpha}(K)$, we may say the sequence exhibits eventual confinement inside the fiber of concept $K$.
@@ 5 Generative Dynamics and Intelligence
Within the Similarity Field Theory framework, this principle is expressed formally. The premise, “given entities exemplifying a concept,” corresponds to a system-state $Z_{p}$ that contains a subset of entities $X_{K}\subseteq X_{p}$, where every entity in $X_{K}$ belongs to a fiber $F_{\alpha}(K)$.
Intelligence with respect to a concept $K$ is the capacity of a generative operator $G$ to produce new entities $E^{\prime}\in G(X_{p})$ that also belong to the fiber $F_{\alpha}(K)$ by satisfying the condition:
@@ 5.2 Creativity as Re-Contextualization
Furthermore, our framework provides a formal perspective on creativity. This is not treated as creation ex nihilo, but as the formation of a new concept initiated by the re-contextualization of an existing entity. This occurs when an entity $E$, already a member of a source concept $K_{\text{source}}$, is discovered to also possess high similarity to a different, nascent target concept $K_{\text{target}}$. Prior to this discovery, the fiber of the target concept, $F_{\beta}(K_{\text{target}})$, might have been empty within the system. The act of recognizing that $S(E,K_{\text{target}})\geq\beta$ populates this fiber for the first time, establishing $K_{\text{target}}$ as a valid concept and positioning $E$ as its foundational instance. The history of WD-40 illustrates this. Its formula ($E$) was an entity in the fiber of “aerospace anti-corrosion agent” ($K_{\text{source}}$). The discovery of its utility in household applications was a recognition of its high similarity to the nascent concept of a “general-purpose lubricant” ($K_{\text{target}}$), making it the first entity to populate that conceptual fiber and create a new market. At an abstract level, creativity can also be characterized as re-contextualization: an entity in the fiber of a source concept is recognized as crossing the threshold for a target concept, thereby making the new fiber non-empty for the first time.
@@ 6 Interpretation in Machine Learning
Similarity Field Theory provides a new geometric lens for understanding typical machine learning models. As discussed earlier in our “Beyond a purely statistical view” subsection, our aim here is not to rehash frequency co-occurrence arguments but to treat networks as knowledge structures of concept fibers; we instantiate this perspective with two canonical cases: neural networks and LLMs [5].
Neural networks as composed similarity fields
Consider a neural network trained to classify images of a cat. Let $K_{\text{Cat}}$ be the concept “cat”. The trained network can be viewed as approximating a similarity field
$S_{\text{NN}}(E_{\text{input}},K_{\text{Cat}})\in[0,1],$
where $E_{\text{input}}$ is an input entity such as a vector of pixel values.
Internal activations are typically unbounded real values. For interpretation, we fix a strictly increasing function $\varphi:\mathbb{R}\to[0,1]$ and define a calibrated activation $a^{\sim}=\varphi(a)\in[0,1]$. This allows us to read a neuron’s response as a similarity value without altering the network’s forward pass. A single neuron in the first hidden layer thus induces a primitive field
$S_{\text{neuron}}(E_{\text{input}},K_{\text{feature},w}):=\varphi(a(E_{\text{input}})),$
where $K_{\text{feature},w}$ denotes the latent feature concept associated with the neuron’s weight vector $w$. For self-similarity we explicitly normalise $S_{\text{neuron}}(K_{\text{feature},w},K_{\text{feature},w})=1$. In practice one can pick a prototype input $E^{\star}$ whose calibrated activation $\varphi(a(E^{\star}))$ is close to the supremum and interpret it as approximately achieving similarity $1$.
At a suitable level of abstraction, the output of a layer (or block) can be viewed as applying a map $g_{\ell}:[0,1]^{m}\to[0,1]$ to a vector of calibrated inputs, where $g_{\ell}$ is coordinatewise monotone and satisfies $g_{\ell}(1,\dots,1)=1$ on chosen prototypes. We do not require each primitive operation inside the layer to be a similarity field by itself; rather, after calibration we interpret the layer-level map as an aggregator over upstream field values. In common architectures this aggregator is implemented by nonnegative weighted sums, convex combinations, or pointwise products of calibrated activations, which keep the range in $[0,1]$ and preserve the prototype value $1$. Iterating this construction over layers yields, at the level of layer outputs, a composed similarity field whose final readout approximates $S_{\text{NN}}(E_{\text{input}},K_{\text{Cat}})$.
Training can then be viewed as similarity maximization: the network parameters are updated so that entities belonging to the cat fiber $F_{\alpha}(K_{\text{Cat}})$ receive higher values of $S_{\text{NN}}(E_{\text{input}},K_{\text{Cat}})$, subject to task constraints. In this view, backpropagation shapes an internal similarity field rather than just fitting an input–output statistic.
@@
Language models as token-level similarity generators
At the language interface, each vocabulary token can be treated as a concept $K_{\text{token}}\in U$, and an input prompt as a finite sequence of such token-concepts whose composition induces a higher-order concept $K_{\text{prompt}}$. The trained language model itself is an entity $K_{\text{Language}}$ that represents a vast meta-concept of language and, simultaneously, a generative operator $G$. For a given token concept $K_{\text{token}}$, its fiber $F_{\alpha}(K_{\text{token}})$ collects entities in $U$ that align with the token’s usage in human language and culture at threshold $\alpha$. When $G$ acts on a prompt, the prediction head compares the synthesized prompt concept $K_{\text{prompt}}$ against every token concept in its vocabulary; the resulting normalized probability distribution over tokens can be interpreted, after calibration, as a monotone surrogate of the similarity values $S(E,K_{\text{prompt}})$. Selecting a high-probability next token $E^{\prime}_{\text{next}}$ therefore approximately enforces $S(E^{\prime}_{\text{next}},K_{\text{prompt}})\geq\alpha$ for some effective threshold $\alpha$, so that the model behaves as a generative operator that tends to remain within the fiber of the prompt concept under the Similarity Field Theory view.
A path towards interpretability
In the SFT view, a trained network is a composition of calibrated similarity fields and can be analysed in terms of the concepts attached to its units. Each neuron computes a value $S_{\text{neuron}}(E_{\text{input}},K_{\text{latent}})$, which we interpret as the similarity between an input entity and a latent concept $K_{\text{latent}}$. In an idealised limit, the concept represented by that neuron is fully captured by its core fiber: the set of all inputs that drive its calibrated activation to the maximum value $1$,
@@
Recovering, even approximately, the inputs in this fiber amounts to recovering what the neuron “means”. From this perspective, interpretability becomes a geometric problem: decomposing the network into its constituent conceptual fibers and understanding how these fibers are composed across layers.
As Similarity Inference